the sides of a triangle measure 18in, 25 in, and 36in. to the nearest degree, what is the largest angle?
I'm seeing a lot of stuff about the law of cosines, but our teacher has never taught us that...I'm in pre-ap geometry

Answer by rothauserc(4718)

The largest angle will be opposite the longest side, let c be the largest side and we use the law of cosines
c^2 = a^2 + b^2 - 2abcos(C)
36^2 = 18^2 + 25^2 - 2*18*25*cos(C)
1296 = 324 + 625 - 900*cos(C)
1296 = 949 - 900*cos(C)
347 = -900*cos(C)
cos(C) = -347/900
cos(C) = -0.385555556
C = 112.678234186
The largest angle = 113 degrees